OpenStudy (anonymous):

Find the limit of (x-2)/(x^2+4x+4) ? When x=-2

OpenStudy (anonymous):

what do you get when you replace \(x\) by \(-2\)?

OpenStudy (anonymous):

-4/0

OpenStudy (anonymous):

wrong.

OpenStudy (anonymous):

Really?

OpenStudy (anonymous):

plug it in.

OpenStudy (anonymous):

you are right, so the limit does not exist

OpenStudy (anonymous):

Misread it sorry.

OpenStudy (anonymous):

DNE?

OpenStudy (anonymous):

4/0

OpenStudy (anonymous):

there is nothing further you can do with \(\frac{-4}{0}\) that is not a number if you had \(\frac{0}{0}\) then you could factor and cancel, and if you had \(\frac{a}{b}\) then that would be your answer yes, it does not exist

OpenStudy (anonymous):

Thank you so much ! (: